OpenStudy (anonymous):

To skate at Roller Heaven, each person must pay a membership fee (m) and a fee for each session (f). Art attended five sessions and paid $25 in all. Later, when he had attended 10 sessions, he paid $40 in all. How much is the membership fee and how much are the sessions. Write a system of two equations with the variables m and f in each to represent this problem.

OpenStudy (anonymous):

m+5f=25 m+10f=40

OpenStudy (anonymous):

m=25-5f subs in eq 2 25-5f+10f=40 5f=15 f=$3 then m=25-5(3) m=10

OpenStudy (anonymous):

f=$3 and m=$10